### v2

# LUSIFER: Language Universal Space Integration for Enhanced Multilingual Embeddings with Large Language Models

Abstract.
Recent advancements in large language models (LLMs) based embedding models have established new state-of-the-art benchmarks for text embedding tasks, particularly in dense vector-based retrieval. However, these models predominantly focus on English, leaving multilingual embedding capabilities largely unexplored. To address this limitation, we present LUSIFER, a novel zero-shot approach that adapts LLM-based embedding models for multilingual tasks without requiring multilingual supervision. LUSIFER’s architecture combines a multilingual encoder, serving as a language-universal learner, with an LLM-based embedding model optimized for embedding-specific tasks. These components are seamlessly integrated through a minimal set of trainable parameters that act as a connector, effectively transferring the multilingual encoder’s language understanding capabilities to the specialized embedding model. Additionally, to comprehensively evaluate multilingual embedding performance, we introduce a new benchmark encompassing 5 primary embedding tasks, 123 diverse datasets, and coverage across 14 languages. Extensive experimental results demonstrate that LUSIFER significantly enhances the multilingual performance across various embedding tasks, particularly for medium and low-resource languages, without requiring explicit multilingual training data. The code and dataset for training are available at: https://github.com/hieum98/lusifer
Large Language Models, Representation Learning, Multilingual Embeddings
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; July 13-18, 2025; Padua, Italyjournalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730029isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Language models
1. Introduction
Text embeddings, which provide dense vector representations of textual content (Mikolov et al., 2013; Devlin et al., 2019), have become fundamental building blocks in modern natural language processing. These embeddings encode semantic information and serve as an important component for numerous downstream applications, ranging from information retrieval and document reranking to classification, clustering, and semantic textual similarity assessment. Recently, the significance of high-quality embeddings has been further amplified by their crucial role in retrieval-augmented generation (RAG) systems (Lewis et al., 2020b). RAG architectures enable large language models (LLMs) to dynamically access and integrate external or proprietary knowledge without the need for model parameter updates, substantially enhancing their adaptability and accuracy (Wang et al., 2023; Liu et al., 2024b; Gao et al., 2024).
The evolution of embedding models has witnessed remarkable advancements, progressing from static word embeddings (Robertson et al., 2009) through contextualized representations (Reimers and Gurevych, 2019; Gao et al., 2021a; Ni et al., 2021a) to state-of-the-art LLM-based embedding models (Wang et al., 2024b) that harness the sophisticated semantic understanding capabilities of large language models. These developments have substantially enhanced performance across various embedding tasks (Luo et al., 2024), achieving unprecedented accuracy in semantic similarity and retrieval applications. However, a critical limitation remains: the predominant focus on English in LLM-based embedding models has created a significant disparity in multilingual capabilities. This gap is especially pronounced in medium and low-resource languages, where English-centric models exhibit substantial performance degradation due to insufficient language-specific training data (Wang et al., 2020; Thakur et al., 2024). While recent advances in multilingual embedding models, particularly those leveraging multilingual pre-trained architectures, have demonstrated promising results in multilingual embedding tasks (Li et al., 2023; Wang et al., 2024c; Chen et al., 2024), their reliance on explicit multilingual supervision for embeddings constrains their applicability primarily to languages with abundant training resources, leaving the challenge of true language-agnostic representation largely unaddressed.
To address this challenge, we present LUSIFER, a novel zero-shot approach that adapts English LLM-based embedding models for multilingual tasks without requiring explicit multilingual supervision. Drawing inspiration from recent advances in multimodal integration (Liu et al., 2024a; Lu et al., 2024), LUSIFER employs a unique architecture that bridges the gap between multilingual understanding and specialized embedding capabilities. At its core, LUSIFER leverages the robust multilingual representations from XLM-R (Conneau et al., 2020) and introduces a learnable connector mechanism to interface with English-optimized LLM embedding models. This approach enables LUSIFER to effectively transfer the multilingual understanding of XLM-R to the target LLM while inheriting advanced embedding capabilities of the LLM. In this way, LUSIFER can achieve effective multilingual representation capabilities without requiring explicit multilingual training data. We conduct comprehensive evaluations of LUSIFER through extensive experiments across 123 diverse datasets spanning 14 languages, focusing on five fundamental embedding tasks: Classification, Clustering, Reranking, Retrieval, and Semantic Textual Similarity (STS). Our experimental results demonstrate that LUSIFER substantially enhances the performance of English-centric LLM-based embedding models, achieving average improvements of 3.19 points across all tasks, with particularly significant gains observed for medium and low-resource languages (up to 22.15 improvement). To validate LUSIFER’s broader applicability and cross-lingual capabilities, we extend our evaluation to cross-lingual tasks using four comprehensive datasets that encompass over 100 languages, including several critically low-resource languages. LUSIFER significantly outperforms existing English-centric embedding models by 5.75 on average in cross-lingual scenarios. These results demonstrate the effectiveness of our approach in enhancing multilingual representation capabilities without explicit multilingual supervision.
The theoretical foundation for LUSIFER’s effectiveness lies in its ability to create a language-agnostic universal space through the integration of a multilingual encoder (Pires et al., 2019; Libovický et al., 2020). We hypothesize that this universal space serves as a bridge between different languages, enabling the target language model to process semantic information independently of the input language. By mapping these language-neutral representations to the target model’s input space, we conjecture that the target LLM can grasp the semantics of these representations, thereby improving the quality of output embeddings across multiple languages. This mechanism allows the model to become less dependent on the specific language of the input, enabling it to better capture semantic information for embedding tasks in languages it rarely encountered during pretraining. Our empirical analysis using t-SNE visualization supports this hypothesis.
Figure 1. Overview of LUSIFER. Left: Align a multilingual encoder with the target English-centric LLM only using English data and a minimal set of trainable parameter. Center: End-to-end representation finetune through contrastive learning on English text-embedding tasks using LoRA. Right: During inference, LUSIFER successfully processes text-embedding tasks across multiple languages.
2. Related Work
2.1. English-centric Embedding Models
Text embedding models have experienced significant advancement in recent years, driven by the evolution of pre-trained language models. Early successes with BERT-based architectures, as demonstrated in Sentence-BERT (Reimers and Gurevych, 2019), SimCSE (Gao et al., 2021a), and DPR (Karpukhin et al., 2020), established the foundation for modern embedding approaches. The field has since progressed to leverage LLMs, with recent works (Wang et al., 2024b; Muennighoff et al., 2024; BehnamGhader et al., 2024; Lee et al., 2024; Man et al., 2024) demonstrating substantial improvements in embedding quality and task performance through the enhanced representational capacity of LLMs (Luo et al., 2024). However, these advances primarily benefit high-resource language applications, as most state-of-the-art LLM-based embedding models are derived from English-centric foundation models (Jiang et al., 2023; Touvron et al., 2023) and trained predominantly on English or high-resource language datasets (Wang et al., 2024b). This bias has resulted in a significant performance gap between high-resource and low-resource languages, limiting the global applicability of these models. Our proposed method, LUSIFER, addresses this limitation by enabling effective multilingual representation without multilingual training data.
2.2. Zero-shot Multilingual Embedding
Multilingual Embedding has evolved through several distinct methodological approaches, each addressing the fundamental challenge of bridging language gaps in embedding tasks. Early successful approaches relied on translation models to enable multilingual understanding (Liu et al., 2020; Shi et al., 2021; Zhang and Misra, 2022). While effective, these methods introduced operational complexity by requiring external translation systems, limiting their practical deployment and scalability. The emergence of multilingual pre-trained language models, particularly XLM-R (Conneau et al., 2020), opened new possibilities for multilingual transfer. Recent works have demonstrated promising results by fine-tuning such models with contrastive learning objectives on multilingual data (Wang et al., 2024c; Chen et al., 2024; Sturua et al., 2024). However, these approaches face two key limitations: they require substantial multilingual training data, and moreover, they do not exploit the sophisticated semantic representations afforded by contemporary English-centric LLM architectures, which have demonstrated superior performance in capturing nuanced semantic relationships.
Recent advances in aligning multilingual and English-centric representations could offer a solution. By combining independently pre-trained representations, a paradigm that has shown remarkable success in multimodal alignment research (Alayrac et al., 2022; Liu et al., 2024a; Lu et al., 2024), these works bridge the gap between visual encoders and language models to enhance visual comprehension. As such, similar principles can be applied to align multilingual representations with LLM-based semantic spaces. While related efforts have explored aligning multiple LLMs for improved reasoning capabilities in multilingual settings (Bansal et al., 2024; Yoon et al., 2024), these approaches primarily target generation tasks and typically require large-scale alignment data. Our work extends these efforts by focusing on embedding tasks and leveraging a minimal set of parameters to align multilingual and English-centric representations, enabling enhanced multilingual representation capabilities without requirement for large-scale multilingual training data.
2.3. Multilingual Embedding Benchmarks
The evaluation landscape for multilingual embedding models has historically been fragmented across various benchmarks, each with significant limitations. While existing benchmarks have made valuable contributions, they often exhibit constrained scope: MINERS (Winata et al., 2024) provides evaluation across multiple languages but is limited to classification and STS tasks with only 11 datasets; XNLI (Conneau et al., 2018), XQuAD (Artetxe et al., 2020), and SIB-200 (Adelani et al., 2024) offer broad language coverage but focus exclusively on classification tasks; and MTEB (Muennighoff et al., 2023), despite its diverse task selection, primarily addresses high-resource languages. To address these limitations, we introduce a comprehensive evaluation framework that encompasses 5 fundamental embedding tasks—Classification, Clustering, Reranking, Retrieval, and STS—across an extensive collection of 123 datasets spanning 14 languages. This holistic approach enables systematic evaluation across both task and language dimensions, providing unprecedented insights into models’ multilingual capabilities. Furthermore, our benchmark extends beyond traditional multilingual evaluation by incorporating cross-lingual tasks, featuring coverage of over 100 languages, including critically low-resource languages that have been historically underrepresented in existing benchmarks. This extensive coverage allows for a more nuanced understanding of embedding models’ performance across the global linguistic landscape.
3. Methodology
Previous works demonstrate that representations of multilingual encoder models exhibit inherent language-agnostic properties, facilitating zero-shot multilingual transfer (Pires et al., 2019; Libovický et al., 2020). Building upon this foundation, we propose LUSIFER, an embedding framework that aligns a multilingual encoder model with a target English-centric LLM’s representational space, enabling the target to encode semantics across multiple languages without extensive multilingual training. This section details our architectural design and two-stage training process for LUSIFER.
3.1. Model Architecture
The core development of LUSIFER lies in its novel approach to enabling multilingual encoding of target LLMs through efficient representation mapping. As illustrated in Figure 1, LUSIFER’s architecture consists of three key components: (1) a multilingual encoder that functions as a language-universal learner, capturing semantic information for diverse languages, (2) a language-agnostic connector that serves as a minimal parametric bridge between representations, and (3) a target LLM optimized for embedding-specific tasks. The multilingual encoder processes input from various languages into a shared semantic space, while the connector, designed with minimal trainable parameters, aligns these universal representations with the target LLM’s native representational space. This alignment enables the target LLM embedding model to effectively leverage multilingual understanding without requiring extensive multilingual training data or architectural modifications.
Following successful approaches in multimodal alignment (Alayrac et al., 2022; Liu et al., 2024a; Lu et al., 2024), we implement the connector as a 2-layers feed-forward network, $\mathbf{FF}$, augmented with a single trainable token appended to the multilingual encoder’s hidden states. Formally, given input tokens $\mathbf{X}_{input}$ (with necessary padding), the multilingual encoder’s hidden states $\mathbf{H}_{enc}$ are transformed to align with the target LLM’s representational space. The resulting aligned hidden states $\mathbf{H}_{align}$ maintain dimensionality compatibility with the target LLM’s hidden states while extending the sequence length by one ($|\mathbf{X}_{input}|+1$): $\mathbf{H}_{align}=[\mathbf{FF}(\mathbf{H}_{enc});\mathbf{t}]$, where $\mathbf{FF}$ is the feed-forward network to align the multilingual encoder’s hidden states with dimension $\mathbf{d}_{e}$ to the target LLM’s hidden states with dimension $\mathbf{d}_{t}$, and $\mathbf{t}\in\mathbb{R}^{d_{t}}$ is the trainable token. Moreover, we employ a masking mechanism to mask any original padding tokens in $\mathbf{H}_{enc}$ to prevent their influence on the target LLM’s processing, ensuring the model focuses on meaningful tokens.
3.2. Training Pipeline
LUSIFER employs a two-stage training process to achieve optimal multilingual representation capabilities. Both stages only require training on English data, leveraging the multilingual encoder’s inherent language-agnostic properties and embedding advantages of LLMs to facilitate zero-shot multilingual transfer.
Figure 2. Overview of tasks and datasets in our benchmark. Crosslingual datasets are marked with a blue shade.
Stage 1: Alignment Training. The initial training stage aligns the multilingual encoder’s representations with the target LLM’s embedding space. Specifically, we optimize the connector parameters $\theta_{c}$ and the multilingual encoder parameters $\theta_{e}$ while keeping the target LLM’s parameters fixed, ensuring stable convergence. The training employs two complementary objectives: (1) A masked reconstruction task where we randomly mask $k\%$ of input tokens such that $\mathbf{X}_{input}=\texttt{mask}(\mathbf{X},k)$, training the model to recover the original sequence $\mathbf{X}_{lm}=\mathbf{X}$. (2) An autoregressive completion task that focuses on next-token prediction, where the model learns to generate the target sequence $\mathbf{X}_{lm}$ conditioned on the input context $\mathbf{X}_{input}$. The training objective for both tasks is formulated as language modeling objective to generate the target sequence $\mathbf{X}_{lm}$ given the input sequence $\mathbf{X}_{input}$. This objective enables local token-level alignment through masked reconstruction task where the model learns to predict the masked tokens by leveraging the context. In addition, it exploits global semantic alignment through autoregressive completion task that encourages the model to capture semantic information of the input sequences to generate the target sequence. As such, our training strategy learns to align the multilingual encoder’s representations with the target LLM’s embedding space while preserving important semantic information of multilingual input sequences. Our training process is conducted using the standard cross-entropy loss function. This stage aims to establish a strong alignment between the multilingual encoder and the target LLM, enabling the target LLM to effectively process multilingual representations produced by the multilingual encoder.
Stage 2: Representation Finetuning. The second stage improves text representations through a contrastive learning process, effectively teaching the model to distinguish between positive and negative examples. Our approach leverages both in-batch negatives sampled from the current training batch and hard-negative examples specifically curated to enhance model training. Additionally, we incorporate bidirectional attention mechanisms within the target LLM, following recent advances in LLM’s representation learning (Muennighoff et al., 2024; BehnamGhader et al., 2024; Lee et al., 2024; Man et al., 2024). This bidirectional context modeling significantly enhances the quality of learned representations by enabling the model to capture both forward and backward dependencies in the input sequence. During this stage, we finetune all components of LUSIFER, including the target LLM, the multilingual encoder, and the connector parameters, to optimize the model’s representation quality for embedding-specific tasks. The goal of this stage is to improve the quality of text representations by leveraging the advanced embedding capabilities of the target LLM while maintaining the multilingual understanding provided by the multilingual encoder.
The two-stage training process enables LUSIFER to effectively align multilingual representations with the target LLM’s embedding space, enhancing the target LLM’s multilingual representation capabilities without requiring explicit multilingual supervision.
4. Experiment
This section presents our experimental methodology and results. We first introduce benchmark datasets and evaluation metrics in Section 4.1, followed by our experimental setup including model implementation, training data, and procedures in Section 4.2. We then present our main findings in Section 4.3 and analyze LUSIFER’s cross-lingual capabilities in Section 4.4. We examine detail performance of LUSIFER across five embedding tasks in Section 4.5 and evaluate LUSIFER’s component effectiveness in Section 4.6. Finally, Section 4.7 visualizes LUSIFER’s representations in multilingual space to demonstrate its language-agnostic capabilities.
Baselines En Es Ru Fr Vi Fa Id Ar Fi Ko Hi Bn Te Sw Avg.
Jina-embeddings-v3* (Sturua et al., 2024) 59.84 61.23 62.88 58.94 66.74 78.35 58.51 64.71 73.57 64.96 64.19 61.54 68.96 49.20 63.83
mGTE-base* (Zhang et al., 2024) 60.40 59.65 61.02 56.20 65.81 73.46 56.55 61.97 68.96 61.22 60.81 58.24 63.58 52.57 61.46
BGE-M3* (Chen et al., 2024) 60.09 60.60 62.37 57.34 70.69 78.97 58.78 64.12 75.60 64.72 64.61 65.31 69.85 54.20 64.80
Multilingual-E5-large* (Wang et al., 2024d) 61.91 61.97 62.91 59.40 71.30 78.08 55.21 63.41 76.53 66.55 63.75 63.67 67.32 51.55 64.54
UDEVER-Bloom-7B* (Zhang et al., 2023) 55.83 56.39 59.73 54.38 64.32 68.70 48.97 55.02 67.60 58.54 55.96 55.13 61.00 47.41 57.78
SimCSE (Gao et al., 2021a) 51.92 51.81 24.90 46.95 31.18 37.12 39.27 29.46 41.64 26.23 25.17 21.54 26.71 38.36 35.16
Contriever (Izacard et al., 2022) 49.29 44.26 26.55 44.05 33.03 39.66 38.33 32.36 45.76 26.47 23.27 22.61 22.64 39.26 34.82
GTE-large (Li et al., 2023) 62.29 51.66 33.49 50.13 38.88 44.67 43.07 30.27 51.98 27.02 20.38 22.97 22.75 41.40 38.64
BGE-en-1.5 (Xiao et al., 2023) 63.27 51.65 32.79 50.84 38.50 49.73 43.28 30.81 51.16 31.11 25.28 26.34 23.02 41.96 39.98
E5-large (Wang et al., 2024a) 60.12 52.41 26.81 51.00 37.99 39.47 43.86 31.32 53.59 28.84 24.57 23.48 22.03 43.25 38.48
ST5-XXL (Ni et al., 2021c) 58.81 60.35 44.42 58.50 41.81 24.66 53.43 25.30 52.46 15.43 18.07 17.10 21.63 38.81 37.91
GTR-XXL (Ni et al., 2021b) 58.12 54.39 41.94 53.21 37.96 24.67 50.08 25.14 53.88 15.23 17.35 15.92 22.12 40.57 36.47
E5-Mistral (Wang et al., 2024b) 66.64 61.84 61.30 59.65 58.58 72.55 58.25 54.43 66.97 62.82 56.23 55.10 47.15 50.61 59.44
LUSIFER (Ours) 57.20 60.14 59.82 59.24 67.69 76.17 59.70 55.60 72.83 65.23 62.37 58.43 69.30 53.12 62.63
Table 1. Comparative analysis of model performance across multiple languages and tasks. The table presents average metrics for each model, with the highest score for each language emphasized in bold. * denotes the models trained on extensive multilingual data.
Baselines MLQARetrieval BelebeleRetrieval STS17 STS22 IndicCrosslingual Avg.
SimCSE (Gao et al., 2021a) 7.41 18.35 39.71 37.95 0.18 20.72
Contriever (Izacard et al., 2022) 9.75 22.94 34.55 41.72 0.03 21.80
GTE-large (Li et al., 2023) 16.99 31.82 37.57 53.79 1.59 28.35
BGE-en-1.5 (Xiao et al., 2023) 16.64 31.19 40.40 50.77 1.11 28.02
E5-large (Wang et al., 2024a) 17.04 31.12 37.90 54.31 1.83 28.44
ST5-XXL (Ni et al., 2021c) 20.82 41.68 56.19 59.02 1.76 35.89
GTR-XXL (Ni et al., 2021b) 20.19 38.02 50.83 60.11 2.74 34.38
E5-Mistral (Wang et al., 2024b) 31.54 54.75 81.12 71.37 21.92 52.14
LUSIFER (Ours) 36.68 57.81 81.09 70.49 43.40 57.89
Table 2. Cross-lingual evaluation results. The table presents average metrics for each model over all languages of the datasets, with the highest score for each language emphasized in bold.
4.1. Benchmark
Figure 2 illustrates the tasks and datasets in our benchmark. Following (Muennighoff et al., 2023), our benchmark includes five fundamental embedding tasks, with the evaluation protocol for each task adapted from the respective original papers. The benchmark involves 123 diverse datasets, including 48 Classification datasets, 24 Clustering datasets, 24 Retrieval datasets, 22 Semantic Textual Similarity STS datasets, and 5 Reranking datasets. The main metrics for each task are as follows: Classification: Accuracy, Clustering: V-measure (Rosenberg and Hirschberg, 2007), Retrieval: nDCG@10, STS: Pearson correlation based on cosine similarity (Reimers et al., 2016), and Reranking: MAP. Following (Lai et al., 2023), our benchmark covers 14 languages including 5 high-resource languages: English (en), Spanish (es), Russian (ru), French (fr), Vietnamese (vi); 6 medium-resource languages: Persian (fa), Indonesian (id), Arabic (ar), Finnish (fi), Korean (ko), Hindi (hi); 3 low-resource languages: Bengali (bn), Telugu (te), Swahili (sw).
Additionally, we evaluate models on cross-lingual retrieval tasks where the models need to perform text embedding tasks with queries and documents in different languages. These tasks feature 5 datasets, including Belebele (Bandarkar et al., 2024), MLQA (Lewis et al., 2020a), STS17, STS22 (Agirre et al., 2016), and IndicCrosslingualSTS (Ramesh et al., 2022), covering over 100 languages, including critically low-resource languages.
(a) Classification tasks
(b) Clustering tasks
Figure 3. Performance comparison of LUSIFER and baseline models on Classification and Clustering tasks.
4.2. Experimental Setup
Implementation Details. LUSIFER encompasses three key components: a multilingual encoder, a connector, and a target LLM. We employ XLM-R-large (Conneau et al., 2020) as the multilingual encoder, Mistral-7B (Jiang et al., 2023) as the English-centric target LLM, and a 2-layer feed-forward network with one trainable token as the connector. To facilitate efficient training, we leverage the LoRA framework (Hu et al., 2022) for training of LUSIFER’s components. Furthermore, we employ GradCache (Gao et al., 2021b), gradient checkpointing, mixed precision training, and FSDP (Zhao et al., 2023) to minimize GPU memory requirements. The LUSIFER architecture and its training code are built on top of the Hugging Face Transformers (Wolf et al., 2020) and Pytorch Lightning libraries (Falcon and team, 2024). We detail the training hyper-parameters for each stage in Table 3
Hyperparameter Alignment Training Representation Finetuning
Batch size 256 256
Learning rate 1.5e-4 5e-5
Learning rate scheduler cosine cosine
Learning rate warm-up ratio 0.1 0.1
Weight decay 0.01 0.01
Grad norm clipping 1.0 1.0
Epochs 2 1
Optimizer AdamW AdamW
Float precision bf16-mixed bf16-mixed
LoRA rank 16 16
LoRA alpha 32 32
Random mask ratio 0.5 -
Number of hardnegatives - 7
Table 3. Training hyperparameters for each stage.
Training Data. We only train LUSIFER on a diverse public English datasets. For alignment training, we use the combination of the English Wikipedia and questions-answering datasets. Specifically, we use subset of Wikitext-103 (Merity et al., 2017) and MSMARCO (Bajaj et al., 2018) for the masked reconstruction and autoregressive completion tasks, respectively. For representation finetuning, we adopt the retrieval datasets as follows: MS MARCO (Bajaj et al., 2018), NQ (Kwiatkowski et al., 2019), PAQ (Lewis et al., 2021), HotpotQA (Yang et al., 2018), SNLI (Bowman et al., 2015), SQuAD (Rajpurkar et al., 2016), ArguAna (Wachsmuth et al., 2018), FiQA (Maia et al., 2018) and FEVER (Thorne et al., 2018). To address the lack of hard negatives in these datasets, we leverage an encoder-based model (Wang et al., 2024a) to select the hard negatives on those datasets. Refer to Table 4 for the number of samples used in each dataset.
Stage Dataset Number of Samples
Alignment Training Wikitext-103 (Merity et al., 2017) 100,000
MSMARCO (Bajaj et al., 2018) 100,000
Representation Finetuning MS MARCO (Bajaj et al., 2018) 100,000
FEVER (Thorne et al., 2018) 100,000
PAQ (Lewis et al., 2021) 100,000
SNLI (Bowman et al., 2015) 100,000
HotpotQA (Yang et al., 2018) 97,800
SQuAD (Rajpurkar et al., 2016) 97,400
FiQA (Maia et al., 2018) 6,420
NQ (Kwiatkowski et al., 2019) 3,420
ArguAna (Wachsmuth et al., 2018) 1,280
Table 4. Number of samples used in each dataset for training. The number of negative samples is included in the total number of samples.
Baselines. We evaluate LUSIFER’s performance across the five fundamental embedding tasks on the benchmark datasets. We make comparisons with a variety of baseline models for embedding tasks which only trained/finetuned on mainly English data. Baselines include the following categories: dense retrieval models with Small Language Model (SLM) backbone: SimCSE (Gao et al., 2021a), Contriever (Izacard et al., 2022), GTE-large (Li et al., 2023), BGE-en-1.5 (Xiao et al., 2023), E5-large (Wang et al., 2024a); and dense retrieval models with Large Language Model (LLM) backbone: GTR-XXL (Ni et al., 2021b), ST5-XXL (Ni et al., 2021c), E5-Mistral (Wang et al., 2024b). Moreover, we include the following state-of-the-art multilingual embedding models which are trained on extensive multilingual data for reference: Jina-embeddings-v3 (Sturua et al., 2024), mGTE-base (Zhang et al., 2024), BGE-M3 (Chen et al., 2024), Multilingual-E5-large (Wang et al., 2024d), and UDEVER-Bloom-7B (Zhang et al., 2023).
4.3. Main Results
Table 1 presents a comprehensive evaluation of LUSIFER’s performance across 14 diverse languages, demonstrating its capabilities in multilingual representation learning. Our model achieves state-of-the-art performance in 10 out of 14 languages, with an average score of 62.63, surpassing the previous benchmark set by E5-Mistral (59.44) by 3.19 points. This improvement is particularly noteworthy given that E5-Mistral utilizes extensive proprietary synthetic data and multilingual training resources. The performance distribution across different language categories reveals LUSIFER’s robust multilingual capabilities. In high-resource languages, while maintaining competitive performance with established benchmarks, our model shows particular strength in medium and low-resource languages. The most striking improvement is observed in Telugu (te), where LUSIFER achieves a remarkable 22.15 points gain over E5-Mistral, underscoring its effectiveness in enhancing representation capabilities for traditionally underrepresented languages. When compared to existing embedding models with SLM backbones, LUSIFER demonstrates substantial improvements over models like E5-large (38.48) and BGE-en-1.5 (39.98) which are trained on English data only, thus further demonstrating the benefits of combining multilingual encoder and LLM’s English-centric for text-embedding tasks in multilingual settings.
Furthermore, LUSIFER achieves competitive performance (62.63) against state-of-the-art multilingual models such as BGE-M3 (64.80) and Multilingual-E5-large (64.54), despite these models requiring extensive multilingual training data. A key advantage of LUSIFER lies in its resource efficiency. While conventional approaches often rely heavily on extensive multilingual training data, our model achieves comparable or superior performance through its innovative alignment mechanism. This efficiency is particularly valuable in scenarios where multilingual training resources are limited or costly to obtain. The model’s ability to generalize effectively across languages while maintaining high performance demonstrates the robustness of our approach.
4.4. Cross-Lingual Evaluation
Figure 4. Performance comparison of LUSIFER and baseline models on Reranking tasks.
(a) Retrieval tasks
(b) STS tasks
Figure 5. Performance comparison of LUSIFER and baseline models on Retrieval and STS tasks.
LUSIFER demonstrates cross-lingual capabilities across multiple languages and tasks, as evidenced by the evaluation results presented in Table 2. Our model achieves a state-of-the-art average score of 57.89, surpassing the previous benchmark set by E5-Mistral (52.14) by a significant margin of 5.75 points. The performance breakdown across different datasets reveals LUSIFER’s consistent superiority. In the MLQARetrieval, LUSIFER achieves a score of 36.68, demonstrating a 5.14 improvement over E5-Mistral’s 31.54. Similarly, on the BelebeleRetrieval dataset, LUSIFER’s score of 57.81 outperforms E5-Mistral’s 54.75, showcasing its robust cross-lingual retrieval capabilities. For semantic similarity tasks, LUSIFER maintains competitive performance, achieving scores of 81.09 and 70.49 on STS17 and STS22 respectively, closely matching or slightly trailing E5-Mistral’s performance. Perhaps the most striking achievement of LUSIFER is its performance on low-resource languages, particularly evident in the IndicCrosslingual dataset. Here, LUSIFER achieves an unprecedented score of 43.40, nearly doubling E5-Mistral’s performance of 21.92. This remarkable improvement demonstrates LUSIFER’s ability to effectively transfer semantic knowledge across language boundaries, particularly benefiting languages with limited resources. These results underscore LUSIFER’s effectiveness in enhancing cross-lingual capabilities through efficient multilingual representation alignment, enabling the model to process text-embedding tasks across multiple languages effectively.
4.5. Task-Specific Performance
Figure 3, 4, 5 present the performance comparison of LUSIFER and baseline models on Classification, Clustering, Reranking, Retrieval, and STS tasks. LUSIFER consistently outperforms the baseline models across 4 out of 5 tasks, with the largest improvements observed in Clustering and Retrieval tasks, especially in the medium and low-resource languages. However, the performance of LUSIFER in the Reranking tasks is slightly worse than the strongest baseline, E5-Mistral model. This discrepancy may be attributed to the task’s complexity and the information loss in the alignment process between the multilingual encoder and the target LLM. Nevertheless, LUSIFER’s strong performance across a variety of tasks and languages highlights its ability to enhance multilingual representations without relying on explicit multilingual training data.
Baselines En Es Ru Fr Vi Fa Id Ar Fi Ko Hi Bn Te Sw Avg.
LUSIFER (Full) 57.20 60.14 59.82 59.24 67.69 76.17 59.70 55.60 72.83 65.23 62.37 58.43 69.30 53.12 62.63
LUSIFER (Connector Only) 35.53 33.98 42.95 33.54 35.68 57.86 35.55 27.60 48.72 34.45 47.57 41.85 46.50 34.66 44.18
LUSIFER (Frozen Multilingual Encoder) 50.99 58.77 58.30 52.73 62.24 75.88 58.11 41.66 70.75 59.53 62.48 55.53 66.24 49.12 58.74
LUSIFER (Alignment Only) 43.32 38.94 45.12 36.75 41.96 64.60 38.38 33.07 52.78 38.08 53.06 47.84 48.34 40.03 44.45
LUSIFER (Representation Finetuning Only) 49.71 58.76 58.08 51.01 62.11 74.01 57.32 40.95 68.47 57.81 59.74 53.53 63.39 47.03 57.28
Table 5. Ablation study results of LUSIFER’s components. The table presents average metrics for each model, with the highest score for each language emphasized in bold.
4.6. Ablation Study
To comprehensively evaluate LUSIFER’s architectural design and training methodology, we conducted an extensive ablation study examining the contribution of each major component. We compared LUSIFER against several ablated variants:
  1. (1)
    Connector-Only: A simplified version using only the finetuning connector during both alignment training and representation finetuning stages, removing the complex interaction between components.
  2. (2)
    Frozen Multilingual Encoder: A variant where the multilingual encoder is freeze throughout training, with only the connector trained during alignment and both connector and target LLM trained during representation finetuning.
  3. (3)
    Alignment Only: The model trained solely with alignment training, omitting the representation finetuning stage to assess the importance of the two-stage training process.
  4. (4)
    Representation Finetuning Only: A version trained using only representation finetuning without initial alignment training.
We evaluated these variants across our standard benchmark suite, measuring performance on cross-lingual transfer, semantic similarity, and downstream task effectiveness. Table 5 presents the comprehensive results of this analysis. The full LUSIFER model achieved the highest average score of 62.63, significantly outperforming all ablated versions. Breaking down the performance impacts: (i) The Connector-Only variant (44.18) showed a 18.45 point, highlighting the importance of allowing the multilingual encoder and the target LLM to be finetuned during training. (ii) The Frozen Multilingual Encoder variant (56.74) performed better than Connector-Only but still fell short by 3.89, demonstrating the value of end-to-end training. (iii) Alignment-Only (44.45) and Representation Finetuning Only (57.28) variants both showed substantial degradation, with drops of 18.18 and 5.35 respectively, indicating the complementary nature of our two-stage training approach, especially the importance of the representation finetuning stage. Through this ablation study, we can attributed that the alignment training stage proves crucial for establishing initial cross-lingual connections, while the representation finetuning stage refines these alignments for specific downstream tasks.
4.7. Model Representation Visualization
(a) E5-Mistral (language)
(b) LUSIFER (language)
Figure 6. t-SNE representation of 200 randomly samples from the SIB200 dataset. The points are colored by the languages.
Figure 6 shows 2D scatter plots of representations from different models for 200 randomly sampled examples from the SIB200 dataset, visualized using t-SNE. The points are colored by the language of the samples. The t-SNE representation of E5-Mistral demonstrates a clearer separation between languages, with distinct clusters for each language. In contrast, the visualization of LUSIFER presents a more mixed distribution of languages, with overlapping clusters across different languages. This observation provides insights into LUSIFER’s lingual-agnostic capabilities, highlighting the model’s ability to bridge the gaps between representation spaces of different languages. These results suggest that LUSIFER’s alignment strategy enables the model to comprehend semantics across multiple languages effectively, facilitating zero-shot multilingual transfer. Overall, our experiments confirm the advantages of the representation alignment strategies in LUSIFER in effectively enabling zero-shot multilingual transfer for LLM-based embedding methods.
5. Conclusion
In this work, we propose LUSIFER, a novel framework that enables effective multilingual representation without explicit multilingual training data. LUSIFER aligns a multilingual encoder with a target English-centric LLM through a minimal set of trainable parameters, facilitating zero-shot multilingual transfer. Our experimental results demonstrate that LUSIFER achieves state-of-the-art performance across diverse languages and tasks, outperforming existing baseline models. Moreover, LUSIFER significantly enhances cross-lingual capabilities, enabling the model to process text-embedding tasks across multiple languages effectively.
Our work provides a promising direction for enhancing multilingual representation capabilities in English-centric embedding models, enabling global applicability without requiring extensive multilingual training data. The alignment strategies employed in LUSIFER ensure that the model can comprehend and process semantic information across different languages, making it a versatile tool for various multilingual applications. Furthermore, the ability of LUSIFER to perform zero-shot multilingual transfer opens up new possibilities for natural language processing tasks in low-resource languages, where obtaining large-scale annotated data is often challenging. By leveraging the strengths of both multilingual encoders and English-centric LLMs, LUSIFER bridges the gap between languages, fostering better multilingual representation learning.
In future work, we plan to explore additional alignment strategies and further investigate the impact of LUSIFER’s components on multilingual representation quality. We also aim to extend our framework to support different modalities such as images and audio and evaluate its performance on a wider range of tasks. Additionally, we will explore the integration of LUSIFER with other state-of-the-art models to further enhance its capabilities. Moreover, we anticipate that LUSIFER will facilitate broader applications of LLM embeddings in downstream tasks, ranging from deep context understanding requirements like sentiment analysis (Gupta et al., 2024) to text style comprehension tasks such as authorship attribution (Rivera-Soto et al., 2021; Man and Nguyen, 2024), thereby contributing to the advancement of natural language processing and information retrieval fields.
Acknowledgements
This research has been supported by the NSF grant # 2239570. This research is also supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract 2022-22072200003. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes, notwithstanding any copyright annotation therein.